OpenStudy (anonymous):

Find sum of the series \[\sum_{n=1}^{\infty}\frac{1}{n^2}\]

OpenStudy (anonymous):

I did integeate geometric series I got this integral \[-\int\limits_{0}^{1}\frac{\ln(1-x)}{x}\]

OpenStudy (anonymous):

ans is (pi)^2/6

OpenStudy (anonymous):

it is geometric series with a=1 and r=1/2 use a/(1-r) to find sum

OpenStudy (anonymous):

http://www2.math.uu.se/~bjorklund/euler.pdf

OpenStudy (anonymous):

Nooo

OpenStudy (anonymous):

why u use integration to solve this

OpenStudy (anonymous):

It's not Geometric, How could it be Geometric?

OpenStudy (anonymous):

Awesome @Ishaan94 how do you find all this stuff

OpenStudy (anonymous):

oh sorry.yes it is not geometric

OpenStudy (anonymous):

Google... since 2004 Lol

OpenStudy (anonymous):

Yeah ,,,,, I mean what do you type in the search box

OpenStudy (anonymous):

Umm I typed 1/n^2 and then suggested searches showed up

OpenStudy (anonymous):

Cool

OpenStudy (anonymous):

It would take me time to digest this thanks

OpenStudy (anonymous):

Me too, thanks for posting the question.

OpenStudy (mr.math):

This can be helpful too http://en.wikipedia.org/wiki/Basel_problem

OpenStudy (anonymous):

We were discussing this thing yesterday (Me, Turing and Badref)

OpenStudy (anonymous):

It took Euler 6 years to get the rigorous proof. :-O

OpenStudy (anonymous):

Btw Robin chapman has this lovely collection of proofs on his homepage : http://empslocal.ex.ac.uk/people/staff/rjchapma/etc/zeta2.pdf

OpenStudy (anonymous):

And here is the url of the epic discussion on M.SE : http://math.stackexchange.com/questions/8337/

OpenStudy (anonymous):

@Ishaan94: And this is how Euler did it http://www.maa.org/editorial/euler/How%20Euler%20Did%20It%2002%20Estimating%20the%20Basel%20Problem.pdf

OpenStudy (anonymous):

Wow

OpenStudy (anonymous):

:-)